### v1

@@ 1 Introduction
To achieve genuine quantum advantage, the development of fault-tolerant quantum computers is required [1]. This poses significant technological challenges, since in practice any quantum computer has to be built from physical platforms, where qubits are actual physical entities that can be manipulated with the goal of performing unitary gate operations [2, 3, 4, 5, 6]. Therefore, it is clear that such operations cannot be dissociated from the underlying physical dynamics, which are inherently noisy [7]. Significant work has been published on how to correct for such gate errors, employing, for instance, sophisticated optimal control theory [8, 9, 10, 11, 12, 13, 14, 15, 16, 17], geometric and dynamical principles [18, 19, 20], or decompositions into “trotterized” or “digitized” gates [21, 22]. In this work, we will take an alternative approach that fully exploits the prowess of counterdiabatic driving.
A set of quantum gates that enables the implementation of any arbitrary $N$-qubit unitary operation is defined as a universal quantum gate set [1]. Naturally, such sets are aimed to be composed of only a small number of gates and low Hilbert space dimensions, to reduce complexity and to ease the controllability of physical systems. One well-known example that forms a universal set for quantum computing comprises all single-qubit gates and the controlled-NOT (CNOT) gate [23, 1], aka “the” Clifford algebra. The CNOT is a two-qubit gate that operates on a target qubit conditioned on the state of the control qubit. In particular, given the regular two-qubit computational basis states $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, it is represented by the unitary matrix,
@@
Thus, the CNOT gate operation maps $|\Psi_{i}\rangle\to|\Psi_{\tau}\rangle=U_{\text{CNOT}}|\Psi_{i}\rangle$, where $|\Psi_{i,\tau}\rangle$ are two states in the computational basis. Notice that $U_{\text{CNOT}}$ does not couple the sectors $\{|00\rangle,|01\rangle\}$ and $\{|10\rangle,|11\rangle\}$, and only in the latter we have the nontrivial operation, $|10\rangle\leftrightarrow|11\rangle$.
@@
For finite driving times $\tau\sim\epsilon^{-1}$, the desired operation deteriorates due to transitions to higher energy states in the spectrum. To correct for these undesired excitations, we introduce an additional control field, the counterdiabatic Hamiltonian [27, 28, 25, 26], which restores the gate operation for arbitrary $\tau$ [29]. Remarkably, even in the presence of environmental noise, we show that the dynamics generated by the bare Hamiltonian $H_{\text{CNOT}}(t)$ plus the counterdiabatic term accurately describes Eq. (1), revealing an interesting and system-independent trade-off between noise strength and driving times.
@@ 2 Hamiltonian gates
A natural way to build Hamiltonian quantum gates from the desired unitary $U$ is by using some inverse engineering method [28, 33, 34, 35, 32, 25]. For unitary dynamics, we can always write $H(t)=i\dot{U}(t)U^{\dagger}(t)$, with $U(t)$ satisfying $U(t)U^{\dagger}(t)=\mathbf{1}$ and $U(0)=\mathbf{1}$ [1]. These two conditions can be used to define a physical $U(t)$, from which the associated Hamiltonian can be obtained directly. For example, following the general algorithm introduced in Ref.[32], it is possible to show that
@@
describes the CNOT operation (1) at time $\tau$, that is,
[…]
where $\mathcal{T}_{>}$ is the time ordering operator (which does not play a role here since $[H(t),H(t^{\prime})]=0$). Moreover, $\varphi(t)$ is an arbitrary function satisfying the boundary conditions $\varphi(0)=2n\pi$ and $\varphi(\tau)=(2n+1)\pi$, with $n$ being an integer, $\sigma^{a=x,y,z}_{j}$ are the Pauli matrices for the qubit $j$, and $\mathbf{1}$ is the $2\times 2$ identity matrix. Notice the non-usual, $ZX$, form of the qubit-qubit interaction in Eq. (2). Interestingly, this form of interaction naturally emerges in some Kitaev quantum spin liquid candidates in two dimensions [36].
Although the Hamiltonian (2) describes the desired CNOT gate (1) in finite time $\tau$, it has trivial dynamics. For example, choosing a linear ramp for the external drive, i.e. $\varphi(t)=\pi t/\tau$, which satisfies the boundary conditions, we obtain a time-independent Hamiltonian. In fact, there are other systematic approaches to the problem of finding time-independent Hamiltonians for quantum gates utilizing machine learning methods [37]. In the present work, we follow a different path and devise driven Hamiltonians that implement quantum gates to highlight the importance of external control for fast gate times. In the next section, we will provide an alternative to Eq. (2), whose ground-state dynamics describes the operation $|10\rangle\leftrightarrow|11\rangle$.
@@ 3 CNOT Hamiltonian
Figure 2: (a) Instantaneous fidelity $\mathcal{F}(t)$, see Eq. (10), between the target state $|11\rangle$ and the evolved state $|\Psi(t)\rangle$ in the adiabatic regime $\tau=200J^{-1}_{1}$. (b) Final fidelity $\mathcal{F}(\tau/2)$ as a function of the driving time $\tau$. (c) Final transition probability $\mathcal{P}(\tau)$ as a function of $\tau$. The blue line represents the numerical solution while the dashed orange line the Landau-Zener formula in Eq. (12). Here we considered $J_{2}(t)=J_{2}t/\tau$. The others parameters are: $J_{2}=10J_{1}$, $g=0.5J_{1}$ and $J_{1}=1$.
@@
$|E_{3}(t)\rangle$ $=$ $|01\rangle,$
$|E_{4}(t)\rangle$ $=$ $|00\rangle,$ (7)
where $\alpha_{\pm}(t)\equiv J_{2}(t)\pm\sqrt{g^{2}+J^{2}_{2}(t)}$. In Fig. 1, we plot the energy spectrum (6) assuming, without loss of generality, linear driving of the second qubit, $J_{2}(t)=J_{2}t/\tau$, with $t\in[-\tau/2,\tau/2]$.
@@
$|E_{1}(-\tau/2)\rangle$ $\simeq$ $|10\rangle,$
$|E_{1}(+\tau/2)\rangle$ $\simeq$ $|11\rangle,$ (8)
@@
$|E_{1}(-\tau/2)\rangle$ $\simeq$ $|11\rangle,$
$|E_{1}(+\tau/2)\rangle$ $\simeq$ $|10\rangle,$ (9)
@@ 4 Solving the dynamics
As a measure of the distance between the quantum state $\rho(t)$ and the desired target state $\rho_{d}\equiv|11\rangle\langle 11|$, we can consider the Uhlmann fidelity [1] $\mathcal{F}(t)=\left(\operatorname{Tr}[\sqrt{\sqrt{\rho_{d}}\rho(t)\sqrt{\rho_{d}}}]\right)^{2}$. Given that both quantum states are pure, we simply have,
$\mathcal{F}(t)$ $=$ $|\langle\Psi(t)|11\rangle|^{2}.$ (10)
Note that $\mathcal{F}(t)$ is bounded between 00 and $1$, with $\mathcal{F}=0$ when there is no overlap between the states, i.e., when they are orthogonal, and $\mathcal{F}=1$ when the two states are identical. Thus, the operation $|10\rangle\leftrightarrow|11\rangle$ is performed successfully if at the end $\mathcal{F}(\tau/2)=1$. In Fig. 2(a) we show $\mathcal{F}(t)$ in the adiabatic regime, $\tau\to\infty$, while in Fig. 2(b) we plot the final fidelity $\mathcal{F}(\tau/2)$ as a function of $\tau$. As anticipated, we can see that the CNOT operation is perfectly realized for very slow driving.
We now return to the decoupled sectors mentioned above. As we can see from Eq. (7), $|\Psi_{i}\rangle$ belongs to $\{|10\rangle,|11\rangle\}$, and therefore $|\Psi(t)\rangle$ does as well. This means that the dynamics occurs in a restricted, two-dimensional Hilbert space and is described by the effective two-level Hamiltonian,
$H_{\text{eff}}(t)$ $=$ $J_{2}(t)\sigma^{z}-g\sigma^{x}-J_{1}\mathbf{1}.$ (11)
Here, $\sigma^{z,x}$ are Pauli matrices in the subspace of $\sigma^{z}_{1}$ projected onto $|1\rangle$. The above Hamiltonian is the Landau-Zener (LZ) Hamiltonian apart from the constant $-J_{1}\mathbf{1}$ [38]. Thus, remarkably, the predictions for the Landau-Zener model are also valid for our case. In particular, the transition probability to $|E_{2}(\tau/2)\rangle$, that is, $\mathcal{P}(\tau)\equiv|\langle\Psi(\tau/2)|E_{2}(\tau/2)\rangle|^{2}$, when $|J_{2}|\gg J_{1},g$, should be given by the LZ formula [38, 39, 40],
@@ 5.1 Finite-time control
where $\dot{H}(t)=\frac{d}{dt}H(t)$ is the reference dynamics (in our case $H(t)\equiv H_{\text{CNOT}}(t)$) and $|E_{n}(t)\rangle$ and $E_{n}(t)$ are the instantaneous eigenstates and eigenenergies of $H(t)$, respectively. The dynamics generated by $H(t)+H_{\text{CD}}(t)$ remains adiabatic for any $\tau$ as long as the system spectrum is gapped during the time evolution, $\Delta(t)\neq 0$. Although it seems that we obtain a general adiabatic speed-up for free, this process becomes highly costly as $\tau\to 0$ [41, 42, 43, 29].
@@
The time-dependent prefactor in the above equation is exactly the same as that appearing in the corresponding $H_{\text{CD}}(t)$ for the Landau-Zener model [44]. Note also that $H_{\text{CD}}(t)$ is entirely spanned in the sector $\{|10\rangle,|11\rangle\}$. This, once more, highlights the fact that we have an effective LZ problem.
@@ 5.2 Noise field
The second term in the above equation, that is, the dissipative term, encodes the stochasticity of the noise process, and $\alpha>0$ measures its strength. Notice that this equation is in the Lindblad form, with only one quantum jump operator given by $L=\sqrt{\alpha}\sigma^{z}_{2}$. Additionally, this same equation also describes the continuous monitoring of $\sigma^{z}_{2}$ at rate $\alpha$ [46].
@@
In Fig. 5, we show the results for the case in which the counterdiabatic Hamiltonian is added. As we can clearly see in both panels, the nonunitary dynamics with $H_{\text{CD}}(t)$ produces excellent results for $\mathcal{F}(\tau/2)$ in short times $\tau<30J^{-1}_{1}$ even for a reasonable noise strength of $\alpha=0.2g$. Therefore, we conclude that the counterdiabatic Hamiltonian (14) can suppress noisy errors in the gate implementation, especially for very short time evolutions.
@@
we will have $\mathcal{F}(\tau/2)\approx 1$ and consequently the desired gate operation will be performed close to perfect. The above relation can be understood again due to the fact that $H_{\text{CD}}(t)\sim\tau^{-1}$. When $H_{\text{CD}}(t)$ is added, the unitary part of Eq. (16) scales as $\tau^{-1}$ [$H_{\text{CNOT}}(t)$ does not scale with $\tau$], while the nonunitary part as $\alpha$. Thus, for larger $\tau$ we need to decrease $\alpha$ in order to suppress the effects of noise, and vice versa. Notice that the relationship $H_{\text{CD}}(t)\sim\dot{H}(t)\sim\tau^{-1}$ is general, i.e., system-independent. Therefore, we expect Eq. (17) to be valid for other Hamiltonian gates described by a master equation of the form (16) in the presence of the corresponding $H_{\text{CD}}(t)$.
@@ 6 $N$-qubits generalization
Once again, we have an effective LZ problem. This presents a significant advantage in terms of the control requirements of the system throughout the desired process.
@@ 7 Concluding remarks
In this work, we have considered the problem of dynamically implementing the CNOT operation $|10\rangle\leftrightarrow|11\rangle$. By locally driving a system of two-interacting qubits, we have shown that the desired gate operation can be performed adiabatically in terms of an effective Landau-Zener (two-level) problem. More realistic situations where errors arise have been analyzed, such as when the gate operation is subjected to non-adiabatic and noise effects. We have shown that counterdiabatic control can properly combat these sources of errors and thus that the control theory based on Eq. (14) is relatively robust against weak noise for short times $\tau$.
Finding new ways to protect unitary gate operations against different forms of errors is essential to quantum computing. Here, we have provided a simple and general way to do this on the basis of shortcuts to adiabaticity. In spirit, our approach is similar to what was proposed in Ref. [41]. However, in our analysis, we have taken an additional step forward and also shown the effectiveness of counterdiabatic control in mitigating errors introduced by possible classical noise in the control fields and the presence of environmental effects.

### v2

@@ 1 Introduction
To achieve genuine quantum advantage, the development of fault-tolerant quantum computers is required [1]. This poses significant technological challenges, since in practice any quantum computer has to be built from physical platforms, where qubits are actual physical entities that can be manipulated with the goal of performing unitary gate operations [2, 3, 4, 5, 6]. Therefore, it is clear that such operations cannot be dissociated from the underlying physical dynamics, which are inherently noisy [7]. Significant work has been published on how to correct for such gate errors, employing, for instance, sophisticated optimal control theory [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], geometric and dynamical principles [20, 21, 22], or decompositions into “trotterized” or “digitized” gates [23, 24, 25]. In this work, we will take an alternative approach that fully exploits the prowess of counterdiabatic driving.
A set of quantum gates that enables the implementation of any arbitrary $N$-qubit unitary operation is defined as a universal quantum gate set [1]. Naturally, such sets are aimed to be composed of only a small number of gates and low Hilbert space dimensions, to reduce complexity and to ease the controllability of physical systems. A standard example of a universal gate set for quantum computing is given by the Hadamard, phase, $\pi/8$, and controlled-NOT (CNOT) gates [26, 1]. The CNOT is a two-qubit gate that operates on a target qubit conditioned on the state of the control qubit. In particular, given the regular two-qubit computational basis states $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$, it is represented by the unitary matrix,
@@
Thus, the CNOT gate operation maps $|\Psi_{i}\rangle\to|\Psi_{\tau}\rangle=U_{\text{CNOT}}|\Psi_{i}\rangle$. Notice that $U_{\text{CNOT}}$ does not couple the sectors $\{|00\rangle,|01\rangle\}$ and $\{|10\rangle,|11\rangle\}$, and only in the latter we have the nontrivial operation, $|10\rangle\leftrightarrow|11\rangle$.
@@
For finite driving times $\tau\sim\epsilon^{-1}$, the desired operation deteriorates due to transitions to higher energy states in the spectrum. To correct for these undesired excitations, we introduce an additional control field, the counterdiabatic Hamiltonian [30, 31, 32, 28, 33, 34, 35, 29, 36], which restores the gate operation for arbitrary $\tau$ [37]. Remarkably, even in the presence of environmental noise, we show that the dynamics generated by the bare Hamiltonian $H_{\text{CNOT}}(t)$ plus the counterdiabatic term accurately describes Eq. (1), revealing an interesting and system-independent trade-off between noise strength and driving times.
@@ 2 Hamiltonian gates
A natural way to build the Hamiltonian $H(t)$ that corresponds to a desired unitary $U$ is by using some inverse engineering method [31, 41, 42, 43, 40, 28]. For unitary dynamics, we can always write $H(t)=i\dot{U}(t)U^{\dagger}(t)$, with $U(t)$ satisfying $U(t)U^{\dagger}(t)=\mathbf{1}$ and $U(0)=\mathbf{1}$ [1]. These two conditions can be used to define a physical $U(t)$, from which the associated Hamiltonian can be obtained directly (in fact, the gauge symmetry $|\tilde{\Psi}(t)\rangle\to V(t)|\Psi(t)\rangle$, for some unitary operator $V(t)$, allows us to build different but equivalent $H(t)$). For example, following the general algorithm introduced in Ref.[40], it is possible to show that
@@
describes the CNOT operation (1) at time $\tau$, that is, $U_{\text{CNOT}}=\mathcal{T}_{>}\exp\left(-i\int_{0}^{\tau}dt\>H(t)\right)$, where $\mathcal{T}_{>}$ is the time ordering operator (which does not play a role here since $H(t),H(t^{\prime})]=0$). Moreover, $\varphi(t)$ is an arbitrary function satisfying the boundary conditions $\varphi(0)=2n\pi$ and $\varphi(\tau)=(2n+1)\pi$, with $n$ being an integer, $\sigma^{a=x,y,z}_{j}$ are the Pauli matrices for the qubit $j$, and $\mathbf{1}$ is the $2\times 2$ identity matrix. Notice the non-usual, $ZX$, form of the qubit-qubit interaction in Eq. (2). Such sort of $ZX$ interaction can be implemented using cross-resonance gates with superconducting qubits [44, 45, 46, 47]. Interestingly, this form of interaction also arises in certain two-dimensional Kitaev quantum spin liquid candidates[48].
Although the Hamiltonian (2) describes the desired CNOT gate (1) in finite time $\tau$, it requires control in the qubit–qubit $ZX$ interaction, which adds both complexity and resource demands relative to simple local single-qubit controls. In addition to this, by choosing a linear ramp $\varphi(t)=\pi t/\tau$, we obtain a trivial time-independent Hamiltonian. In the present work, we follow a different path and devise driven Hamiltonians that implement quantum gates to highlight the importance of external local control for fast gate times. In the next section, we will provide an alternative to Eq. (2), whose ground-state dynamics describes the operation $|10\rangle\leftrightarrow|11\rangle$.
@@ 3 CNOT Hamiltonian
Figure 2: (a) Instantaneous fidelity, $\mathcal{F}(t)=|\langle\Psi(t)|11\rangle|^{2}$ with $|\Psi(0)\rangle=|E_{1}(-\tau/2)\rangle$, between the target state $|11\rangle$ and the evolved state $|\Psi(t)\rangle$ in the adiabatic regime $\tau=200J^{-1}_{1}$. (b) Final fidelity $\mathcal{F}(\tau/2)$ as a function of the driving time $\tau$. (c) Final transition probability $\mathcal{P}(\tau)$ as a function of $\tau$. The blue line represents the numerical solution while the dashed orange line the Landau-Zener formula in Eq. (9). Here we considered $J_{2}(t)=J_{2}t/\tau$. The others parameters are: $J_{2}=10J_{1}$, $g=0.5J_{1}$ and $J_{1}=1$.
@@
$|E_{3}(t)\rangle$ $=$ $|01\rangle,\>\>\>\>\>\>|E_{4}(t)\rangle\>\>=\>\>|00\rangle,$ (6)
where $\alpha_{\pm}(t)\equiv J_{2}(t)\pm\sqrt{g^{2}+J^{2}_{2}(t)}$. In Fig. 1, we plot the energy spectrum (5) assuming, without loss of generality, linear driving of the second qubit, $J_{2}(t)=J_{2}t/\tau$, with $t\in[-\tau/2,\tau/2]$ (negative times here mean the direction of the applied field $J_{2}(t)$). The crossings seen in Fig. 1(a) occur between energy levels from the decoupled sectors. As a result, for the gate operation $|10\rangle\leftrightarrow|11\rangle$, such crossings do not exist, as explicitly shown in Fig. 1(b).
@@
$|E_{1}(-\tau/2)\rangle$ $\simeq$ $|10\rangle,\>\>\>\>|E_{1}(+\tau/2)\rangle\>\>\>\simeq\>\>\>|11\rangle,$ (7)
@@
$|E_{1}(-\tau/2)\rangle$ $\simeq$ $|11\rangle,\>\>\>\>|E_{1}(+\tau/2)\rangle\>\>\>\simeq\>\>\>|10\rangle,$ (8)
@@ 4 Solving the dynamics
As a measure of the distance between the quantum state $\rho(t)$ and the desired target state $\rho_{d}\equiv|11\rangle\langle 11|$, we can consider the Uhlmann fidelity [1] $\mathcal{F}(t)=\left(\operatorname{Tr}[\sqrt{\sqrt{\rho_{d}}\rho(t)\sqrt{\rho_{d}}}]\right)^{2}$. Given that both quantum states are pure, we simply have $\mathcal{F}(t)=|\langle\Psi(t)|11\rangle|^{2}$. In Fig. 2(a) we show $\mathcal{F}(t)$ in the adiabatic regime, $\tau\to\infty$, while in Fig. 2(b) we plot the final fidelity $\mathcal{F}(\tau/2)$ as a function of $\tau$. As anticipated, we can see that the CNOT operation is perfectly realized for very slow driving.
We now return to the decoupled sectors mentioned above. As we can see from Eq. (6), $|\Psi_{i}\rangle$ belongs to $\{|10\rangle,|11\rangle\}$, and therefore $|\Psi(t)\rangle$ does as well. This means that the dynamics occurs in a restricted, two-dimensional Hilbert space $\{|10\rangle,|11\rangle\}$ and is described by the effective two-level Hamiltonian $H_{\text{eff}}(t)=J_{2}(t)\sigma^{z}-g\sigma^{x}-J_{1}\mathbf{1}$. Here, $\sigma^{z,x}$ stand for the Pauli matrices acting in this subspace. The above Hamiltonian is the Landau-Zener (LZ) Hamiltonian apart from the constant $J_{1}\mathbf{1}$ [49]. Thus, remarkably, the predictions for the Landau-Zener model are also valid for our case. In particular, the transition probability to $|E_{2}(\tau/2)\rangle$, that is, $\mathcal{P}(\tau)\equiv|\langle\Psi(\tau/2)|E_{2}(\tau/2)\rangle|^{2}$, when $|J_{2}|\gg J_{1},g$, should be given by the LZ formula [49, 50, 51],
@@ 5.1 Finite-time control
where $\dot{H}(t)=\frac{d}{dt}H(t)$ is the reference dynamics (in our case $H(t)\equiv H_{\text{CNOT}}(t)$) and $|E_{n}(t)\rangle$ and $E_{n}(t)$ are the instantaneous eigenstates and eigenenergies of $H(t)$, respectively. The dynamics generated by $H(t)+H_{\text{CD}}(t)$ remains adiabatic for any $\tau$ as long as the system spectrum is gapped during the time evolution, $\Delta(t)\neq 0$. Although it seems that we obtain a general adiabatic speed-up for free, this process becomes highly costly as $\tau\to 0$ [52, 53, 54, 55, 37]. Precisely, it was shown in [52] that for a two-qubit problem the cost scales as $\tau^{-1}$ when $\tau\to 0$.
@@
The time-dependent prefactor in the above equation is exactly the same as that appearing in the corresponding $H_{\text{CD}}(t)$ for the LZ model [56]. Note also that $H_{\text{CD}}(t)$ is entirely spanned in the sector $\{|10\rangle,|11\rangle\}$. This, once more, highlights the fact that we have an effective LZ problem. In Refs. [57, 54, 58], the counterdiabatic dynamics of the LZ model was experimentally realized.
@@ 5.2 Noise field
[…]
@@
In Fig. 5, we show the results for the case in which the counterdiabatic Hamiltonian is added. As we can clearly see in both panels, the nonunitary dynamics with $H_{\text{CD}}(t)$ produces excellent results for $\mathcal{F}(\tau/2)$ in short times $\tau<30J^{-1}_{1}$ even for a reasonable noise strength of $\alpha=0.2g$. Therefore, we conclude that the counterdiabatic Hamiltonian (11) can avoid noisy errors in the gate implementation, especially for very short time evolutions.
@@
we will have $\mathcal{F}(\tau/2)\approx 1$ and consequently the desired gate operation will be performed close to perfect. The above relation can be understood again due to the fact that $H_{\text{CD}}(t)\sim\tau^{-1}$. When $H_{\text{CD}}(t)$ is added, the unitary part of Eq. (13) scales as $\tau^{-1}$ $H_{\text{CNOT}}(t)$ does not scale with $\tau$], while the nonunitary part as $\alpha$. Thus, for larger $\tau$ we need to decrease $\alpha$ in order to counteract the effects of noise, and vice versa. Notice that the relationship $H_{\text{CD}}(t)\sim\dot{H}(t)\sim\tau^{-1}$ is general, i.e., system-independent. Therefore, we expect Eq. (14) to be valid for other Hamiltonian gates described by a master equation of the form (13) in the presence of the corresponding $H_{\text{CD}}(t)$.
@@ 6 $N$-qubits generalization
Once again, we have an effective LZ problem. This represents a significant advantage in designing the qubit–qubit couplings necessary to implement the gate operation $|11\cdots 10\rangle\leftrightarrow|11\cdots 11\rangle$ in finite time [52].
@@ 7 Concluding remarks
In this work, we have considered the problem of dynamically implementing the CNOT operation $|10\rangle\leftrightarrow|11\rangle$. By locally driving a system of two-interacting qubits, we have shown that the desired gate operation can be performed adiabatically in terms of an effective Landau-Zener (two-level) problem. More realistic situations where errors arise have been analyzed, such as when the gate operation is subjected to non-adiabatic and noise effects. We have shown that counterdiabatic control can properly avoid these sources of errors and thus that the control theory based on Eq. (11) is relatively robust against weak noise for short times $\tau$.
Finding new ways to protect unitary gate operations against different forms of errors is essential to quantum computing. Here, we have provided a simple and general way to do this on the basis of shortcuts to adiabaticity. In spirit, our approach is similar to what was proposed in Ref. [52]. However, in our analysis, we have taken an additional step forward and also shown the effectiveness of counterdiabatic control in mitigating errors introduced by possible classical noise in the control fields and the presence of environmental effects. As a future work, it would be interesting to make a comprehensive comparative analysis of the performance of different quantum control methods [61] in achieving noise robust quantum gates.